# SpectrAx: Spectral Search of Axion-Like Particles Using Multi-Band Observations of Galaxy Clusters from SKA, SO, CMB-S4 and eROSITA

### v1

@@ 4.4 Axion signal in CMB map
The ALP signal can be probed from polarization measurements of the CMB along the cluster line of sight. As explained in Sec. 4.4, the ALP distortion signal will be visible as a disk, with its size depending on the mass of the ALPs being formed [38, 39]. These fluctuations in the cluster will increase with frequency in the microwave spectrum, while the regions surrounding the cluster regions will bear the smooth CMB characteristics, with fluctuations that will be independent of frequency. We show the small-scale anisotropies around a galaxy cluster in different frequency channels accessible from the upcoming CMB missions in Fig. 8.
@@
To be able to probe the ALP signal disk, it is necessary that the cluster be resolvable in the CMB. Also, the beam size of the instrument should be smaller than the coherence length of the magnetic field, so that depolarization of the signal does not occur. We expect that to be the case with the upcoming high-resolution experiments like SO, CMB-S4, etc. Also, the scale of variation of the electron density should be smaller than the scale of the ordered magnetic field. This will let us neglect the effects of turbulence in the magnetic field, which otherwise would lead to partial depolarization of the signal.
@@
Figure 8: This figure depicts the change that is caused in the polarization fluctuations of the CMB at small angular scales (HEALPix NSIDE $=$ 4096) corresponding to the region of $\sim 37\times 37$ sq. arcminutes around a cluster in the absence (top left) and presence of ALP signal ($g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$ and $m_{a}=10^{-14}\,\mathrm{eV}$). These fluctuations increase with frequency in the microwave spectrum. The circular region that the cluster occupies shows an increase in fluctuations, while the smooth CMB characteristics are visible in the surrounding region.
@@
For our mock analysis, we have considered ALPs of masses $10^{-14}$ eV, with a tolerance of ±$\pm 15\%$ to be forming in the clusters with a uniform coupling constant of $5\times 10^{-12}\,\mathrm{GeV^{-1}}$ in the mock data. We consider a uniformly random transverse magnetic field direction. To take this into account, the signal is generated for various clusters at different frequencies assuming a uniformly random polarization direction. The case for zero coupling is considered in Appendix A and highlights that the method is unbiased.
@@
The detectability of the ALP signal is still affected by various contaminants. The CMB is contaminated by foregrounds, such as the galactic synchrotron emission and thermal dust emission. These are highly dominating in the galactic plane and need to be cleaned or accounted for in the analysis. For our mock analysis, we generate the ALP signal in the clusters, obtain the CMB polarization maps from CAMB, and the foregrounds galactic dust and synchrotron emission from PySM [101]. We use the dust model "d-3", with a spatially varying spectral index. Also, we use the synchrotron model "s-3" which has a curved spectrum, steepening or flattening with frequency. The synchrotron emission is strong at frequencies less than $\nu=70$ GHz, while thermal dust emission impacts the signal for frequencies greater than $=200$ GHz. Masking the galactic plane to mitigate their impact thus becomes important, although at a cost of the sky fraction available for analysis. We do not consider the polarized SZ signals in our analysis as these are weak (of the order of nano-Kelvins) and can be separated based on the distinct spectral and spatial variation of the ALP signal [39, 102, 103, 104, 105, 106, 80, 107, 108, 78].
@@ 4.5 Cleaning of the foregrounds
4.6 Multi-band Bayesian framework to search for axions
@@
The coupling constant determination involves the multi-band hierarchical Bayesian analysis. This technique involves marginalizing the nuisance parameters to obtain posteriors for required parameters. The posteriors for the magnetic field and electron density profile parameters are marginalized over to obtain the posterior on the coupling constant $g_{a\gamma}$ using the ILC cleaned map. The error in redshift is obtained from the optical surveys. The errors on the magnetic field, electron density, and redshifts get propagated to provide constraints on the coupling constant $g_{a\gamma}$. A uniform prior on $g_{a\gamma}$ is used for values from $\mathrm{10^{-14}}$ to $\mathrm{1.5\times 10^{-11}\,GeV^{-1}}$ in this analysis. This is given as a $p+q+1$ order integration, with n being the number of magnetic field parameters, q being the number of electron density parameters that are being marginalized over and z, the redshift distribution
@@
In the absence of an ALPs signal, the terms associated with $Q_{\mathrm{ax}}$ and $U_{\mathrm{ax}}$ are zero. However, in the presence of the ALPs signal, the second term in Equ.(4.15) need not vanish after summing over many pixels, if the number of pixels are less. So, we need to calculate this number for an accurate estimation of the coupling constant. For a pessimistic choice of $g_{a\gamma}=10^{-11}$ GeV-1, close to the current bound from CAST [114], we calculate the covariance matrix. If we consider coupling strength to be zero in the covariance matrix, then the uncertainty in the inference from the CMB map would reduce, but could also lead to biased results if ALPs exist with a non-zero coupling constant. In the SpectrAx code, the covariance matrix can be calculated for both zero and non-zero values of the coupling constant. The covariance will be estimated using different realizations of the ALP signal, calculated with a high value of the coupling constant ($10^{-11}\,\mathrm{GeV^{-1}}$ in this analysis) and varying electron density and magnetic field profiles. The use of a high coupling constant will give us conservative constraints on the coupling constant, as it increases the covariance. The covariance for a cluster region due to various components (CMB, ALP signal, foregrounds, instrument noise, etc.) is obtained as the ensemble average of the variance of multiple cluster-sized regions (with different profiles and ALP signals for a high coupling constant) in the ILC cleaned map.
@@ 5 Results
We present the constraints obtained on the coupling constant using the hierarchical Bayesian framework for different sets of multi-band surveys. We have the magnetic field constraints from our mock analysis using SKA configuration, electron density constraints from eROSITA configuration, and CMB observations from SO and CMB-S4.
Figure 9: Constraints on $g_{a\gamma}$ for a combination of SKA, eROSITA, and SO for the 10 clusters. The bold black line represents the merged posterior. True value is shown as a dashed vertical line, which is $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV}^{-1}$.
Figure 10: Constraints on $g_{a\gamma}$ for a combination of SKA, eROSITA and CMB-S4 for the 10 clusters shown in different colours. The bold black line represents the merged posterior. True value is shown as a dashed vertical line, which is $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV}^{-1}$.
SpectrAx gives an accurate inference of the ALPs coupling strength for various ALP masses. We show the constraints using a combination of SKA $+$ eROSITA $+$ SO in Fig. 9 and SKA $+$ eROSITA $+$ CMB-S4 in Fig. 10. The various lines represent constraints from different clusters, with the bold black line in each plot obtained by merging constraints from all clusters. The shaded region represents the acceptable values for the ALP coupling constant. It can provide constraints on the coupling constant in $\sim 50$ core-hours per 1000 pixels.
SpectrAx gives reliable results in the case of no coupling ($g_{a\gamma}=0$) as has been shown in Appendix A. This case will give an upper bound on the photon-ALP coupling constant, and can only be further probed with higher resolution and sensitive detectors.
@@ 6 Conclusion
If galaxy clusters are resolvable in different EM bands, SpectrAx can infer their magnetic field from radio synchrotron emission, and electron density and temperature from X-ray emission. Using redshift measurements from optical surveys 4.1, SpectrAx will be able to obtain constraints on the ALP coupling constant $g_{a\gamma}$ under a hierarchical Bayesian framework through the CMB polarization observations from detectors like SO and CMB-S4 (see Sec. 4.6). SpectrAx is able to infer the ALP coupling constant using 6 nuisance parameters in $\approx 50$ core-hours per 1000 pixels (see Table 6). All the values quoted above are model and parameters dependent, but SpectrAx performs at an optimum pace and efficiency.
@@
In the near future, the number of galaxy clusters detectable with the ongoing/upcoming surveys in radio, CMB, optical, and X-ray bands, will be a few tens of thousands with much higher angular resolutions than now. SpectrAx can constrain radial profiles of magnetic field and electron density using these observations, along with finding beyond standard model physics. The method developed in this framework will pave the way for astrophysical and cosmological studies in this era of precision cosmology in the coming years.
Appendix A Constraints for zero coupling ($g_{a\gamma}=0$)
The results for a non-zero value of the coupling constant are presented in Sec. 5. Here we check the reliability of SpectrAx in case of non-existence of ALPs ($g_{a\gamma}=0$). We present the results of zero coupling for the clusters in Fig. 11. These constraints are obtained using a combination of SKA $+$ eROSITA $+$ CMB-S4. The constraints for these clusters using the same configurations for a coupling constant of $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$ can be seen in Fig. 10.
Currently the bounds on coupling constant from CAST are at $g_{a\gamma}=6.6\times 10^{-11}\,\mathrm{GeV}^{-1}$. The upcoming detectors will be observing tens of thousands of galaxy clusters. Constraints from thousands of clusters will provide much better bounds on the ALP coupling constant. Similar analysis at the power spectrum level can also be performed for resolved as well as unresolved galaxy clusters [39, 40].

### v3

@@ 4.4 ALP signal in CMB map
The ALP signal can be probed from polarization measurements of the CMB along the cluster line of sight. As explained in Sec. 3.5, the ALP distortion signal for a resolved cluster will be visible as a disk, with its size depending on the mass of the ALPs being formed [38, 39]. These fluctuations in the cluster will increase with frequency in the microwave spectrum, while the regions surrounding the cluster regions will bear the smooth CMB characteristics, with fluctuations that will be independent of frequency. We show the small-scale anisotropies around a resolved galaxy cluster in different frequency channels accessible from the upcoming CMB missions in Fig. 8.
@@
To be able to probe the ALP signal disk, it is necessary that the cluster be resolvable in the CMB. Also, the beam size of the instrument should be smaller than the coherence length of the magnetic field, so that depolarization of the signal does not occur. We expect that to be the case with the upcoming high-resolution experiments like SO, CMB-S4, etc. Also, the scale of variation of the electron density should be smaller than the scale of the ordered magnetic field. This will let us neglect the effects of turbulence in the magnetic field, which otherwise would lead to partial depolarization of the signal. If multiple mass ALPs are formed, there will be line of sight depolarization depending on the orientation of magnetic field at the conversion locations.
@@
Figure 8: This figure depicts the change that is caused in the polarization fluctuations of the CMB at small angular scales (HEALPix NSIDE $=$ 4096) corresponding to the region of $\sim 44\times 44$ sq. arcminutes around a resolved cluster in the absence (top left, with CMB being independent of frequency) and presence of ALP signal ($g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$ and $m_{a}=10^{-14}\,\mathrm{eV}$). These fluctuations increase with frequency in the microwave spectrum. The circular region that the cluster occupies shows an increase in fluctuations, while the smooth CMB characteristics are visible in the surrounding region.
@@
For our mock analysis, we have considered ALPs of masses $10^{-14}$ eV, with a tolerance of ±$\pm 15\%$ to be forming in the clusters with a uniform coupling constant of $5\times 10^{-12}\,\mathrm{GeV^{-1}}$ in the mock data. We consider a uniformly random transverse magnetic field direction. To take this into account, the signal is generated for various clusters at different frequencies assuming a uniformly random polarization direction. Also, the case for zero coupling is considered in Sec. 5 and highlights that the method is unbiased.
@@
The detectability of the ALP signal is still affected by various contaminants. The CMB is contaminated by foregrounds, such as the galactic synchrotron emission and thermal dust emission. These are highly dominating in the galactic plane and need to be cleaned or accounted for in the analysis. For our mock analysis, we generate the ALP signal in the clusters, obtain the CMB polarization power spectrums from CAMB and the corresponding maps using the synfast routine in HealPy [122], and the foregrounds galactic dust and synchrotron emission from PySM [123]. We use the dust model "d-3", with a spatially varying spectral index. Also, we use the synchrotron model "s-3" which has a curved spectrum, steepening or flattening with frequency. The synchrotron emission is strong at frequencies less than $\nu=70$ GHz, while thermal dust emission impacts the signal for frequencies greater than $=200$ GHz. Masking the galactic plane to mitigate their impact thus becomes important, although at a cost of the sky fraction available for analysis. We do not consider the polarized SZ signals in our analysis as these are weak (of the order of nano-Kelvins) and can be separated based on the distinct spectral and spatial variation of the ALP signal [39, 124, 125, 126, 127, 128, 80, 129, 130, 78].
@@ 4.5 Cleaning of the foregrounds
4.6 Multi-band Bayesian framework to search for ALPs
@@
The coupling constant determination involves the multi-band hierarchical Bayesian analysis. This technique involves marginalizing the nuisance parameters to obtain posteriors for required parameters. The posteriors for the magnetic field and electron density profile parameters are marginalized over to obtain the posterior on the coupling constant $g_{a\gamma}$ using the ILC cleaned map. The error in redshift is obtained from the optical surveys. The errors on the magnetic field, electron density, and redshifts get propagated to provide constraints on the coupling constant $g_{a\gamma}$. A uniform prior on $g_{a\gamma}$ is used for values from $\mathrm{10^{-14}}$ to $\mathrm{1.5\times 10^{-11}\,GeV^{-1}}$ in this analysis, with an increased range from Table 2, to check for the stability of the results in case of variation in prior range. This is given as a $p+q+1$ order integration, with $p$ being the number of magnetic field parameters, $q$ being the number of electron density parameters that are being marginalized over and $z$, the redshift distribution
@@
In the absence of an ALPs signal, the terms associated with $Q_{\mathrm{ax}}$ and $U_{\mathrm{ax}}$ are zero. However, in the presence of the ALPs signal, the second term in Equ.(4.15) need not vanish after summing over many pixels, if the number of pixels is less. So, we need to calculate this number for an accurate estimation of the coupling constant. For a pessimistic choice of $g_{a\gamma}=10^{-11}$ GeV-1, close to the current bound from CAST [109], we calculate the covariance matrix. If we consider coupling strength to be zero in the covariance matrix, then the uncertainty in the inference from the CMB map would reduce, but could also lead to biased results if ALPs exist with a non-zero coupling constant. In the SpectrAx code, the covariance matrix can be calculated for both zero and non-zero values of the coupling constant. The covariance will be estimated using different realizations of the ALP signal, calculated with a high value of the coupling constant ($10^{-11}\,\mathrm{GeV^{-1}}$ in this analysis) and varying electron density and magnetic field profiles. The use of a high coupling constant will give us conservative constraints on the coupling constant, as it increases the covariance. The covariance for a cluster region due to various components (CMB, ALP signal, foregrounds, instrument noise, etc.) is obtained as the ensemble average of the variance of multiple cluster-sized regions (with different profiles and ALP signals for a high coupling constant) in the ILC cleaned map.
@@ 5 Results
We present the constraints obtained on the coupling constant using the hierarchical Bayesian framework for different sets of multi-band surveys for two scenarios of coupling strengths, (i) $g_{a\gamma}=0$ and (ii) $g_{a\gamma}=5\times 10^{-12}$ GeV-1. We have the magnetic field constraints from our mock analysis using SKA configuration, electron density constraints from eROSITA configuration, and CMB observations from SO and CMB-S4.
Figure 9: Constraints on $g_{a\gamma}$ ($m_{a}=10^{-14}$ eV) for a combination of SKA, eROSITA, and SO for the 10 clusters. The bold black line represents the merged posterior. True value is shown as a dashed vertical line, which is $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV}^{-1}$.
Figure 10: Constraints on $g_{a\gamma}$ ($m_{a}=10^{-14}$ eV) for a combination of SKA, eROSITA and CMB-S4 for the 10 clusters shown in different colours. The bold black line represents the merged posterior. True value is shown as a dashed vertical line, which is $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV}^{-1}$.
5.1 Zero ALP coupling ($g_{a\gamma}=0$):
SpectrAx gives reliable results in the case of no coupling ($g_{a\gamma}=0$) as well. This case will give an upper bound on the photon-ALP coupling constant, and can only be further probed with higher resolution and sensitive detectors.
Here we check the reliability of SpectrAx in case of non-existence of ALPs ($g_{a\gamma}=0$). We present the results of zero coupling for the clusters in Fig. 11. These constraints are obtained using a combination of SKA $+$ eROSITA $+$ CMB-S4. The constraints for these clusters using the same configurations for a coupling constant of $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$ can be seen in Fig. 10.
Currently the bounds on coupling constant from CAST are at $g_{a\gamma}=6.6\times 10^{-11}\,\mathrm{GeV}^{-1}$. The upcoming detectors will be observing tens of thousands of galaxy clusters. Constraints from thousands of clusters will provide much better bounds on the ALP coupling constant. Similar analysis at the power spectrum level can also be performed for resolved as well as unresolved galaxy clusters. The bounds that can be obtained from a large number of clusters accessible with SO and CMB-S4 for the case of no-ALPs are studied in [39] (for the case of $10^{-13}$ eV mass ALPs) and [40]. We are able to obtain bounds on the ALP coupling constant, up to 95% confidence interval for ALP mass $10^{-13}$ eV at $g_{a\gamma}<5.2\times 10^{-12}\,\mathrm{GeV^{-1}}$ using clusters up to redshift $z=1$ with SO (which will observe 24000 clusters upto $z\sim 2$), and $g_{a\gamma}<3.6\times 10^{-12}\,\mathrm{GeV^{-1}}$ with CMB-S4 (which will observe 70000 clusters up to $z\sim 2$).
5.2 Non-zero ALP coupling ($g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$)
SpectrAx gives an accurate inference of the ALPs coupling strength for various ALP masses. Here we analyze the case for $10^{-14}$ eV ALP mass. We show the constraints using a combination of SKA $+$ eROSITA $+$ SO in Fig. 9 and SKA $+$ eROSITA $+$ CMB-S4 in Fig. 10. The various lines represent constraints from different clusters, with the bold black line in each plot obtained by merging constraints from all clusters, by multiplying their posteriors. The shaded region represents the acceptable values for the ALP coupling constant. It can provide constraints on the coupling constant in $\sim 50$ core-hours per 1000 pixels.
@@ 6 Conclusion
If galaxy clusters are resolvable in different EM bands, SpectrAx can infer their magnetic field from radio synchrotron emission, and electron density and temperature from X-ray emission. Using redshift measurements from optical surveys 4.1, SpectrAx will be able to obtain constraints on the ALP coupling constant $g_{a\gamma}$ under a hierarchical Bayesian framework through the CMB polarization observations from detectors like SO and CMB-S4 (see Sec. 4.6). SpectrAx is able to infer the ALP coupling constant using 6 nuisance parameters in $\approx 50$ core-hours per 1000 pixels (see Table 6). SpectrAx uses parallelization in a way that can constrain the parameters for multiple galaxy clusters together, thus data from multiple clusters can be analyzed alongside. This enables an efficient use of multiple core machines and saves time. All the values quoted above are model and parameters dependent, but SpectrAx performs at an optimum pace and efficiency in being able to constrain complex models.
@@
In the near future, the number of galaxy clusters detectable with the ongoing/upcoming surveys in radio, CMB, optical, and X-ray bands, will be a few tens of thousands with much higher angular resolutions than now. SpectrAx can constrain radial profiles of magnetic field and electron density using these observations, along with finding beyond standard model physics. The pipeline will be made publicly available for ALPs analysis from CMB data after its demonstration on data from CMB experiments such as the SPTpol [135] and ACTpol [136, 137]. The method developed in this framework will pave the way for astrophysical and cosmological studies in this era of precision cosmology in the coming years.
Appendix A Convergence of MCMC chains for the inference of ALPs coupling constant
Figure 12: We show five randomly chosen walkers (out of 25 walkers) for the ALPs coupling constant $g_{a\gamma}$ as a function of the MCMC steps obtained using SpectrAx for a single galaxy cluster.
We show the convergence of MCMC chains for one of the galaxy clusters in Fig. 12. Here we have considered 25 walkers (in the figure we show 5 of them), which are represented by different colours. The Gelman-Rubin statistic for this case is 0.99 . The variation in the estimation of nuisance magnetic field and electron density parameters (which are marginalized over) leads to the high deviations from the true value when the number of steps is low. The chains converge to the true value of $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$ when the number of steps increases, accompanied with a decrease in departure of the walkers from the true value. Thus, the earlier part of the chains needs to be discarded. Moreover, thinning is required in the rest of the chain to keep the samples independent. The thinning is performed for all clusters at about 20 steps after discarding 18000 steps. We perform these operations to obtain the posteriors on the ALP coupling constant from different clusters as shown in Sec. 5.